OpenStudy (livya15):

Bob ha Regular tickets are $4 and Premium Tickets are $6. Bob sold 450 tickets and the amount of money was $2340. HOw many tickets were sold?

OpenStudy (anonymous):

Set up a system..... here's the two equations: x+y=540 4x+6y=2340

OpenStudy (anonymous):

then solve it

OpenStudy (anonymous):

I'll show you how, using elimination

OpenStudy (anonymous):

I mean substitution sorry

OpenStudy (anonymous):

So x+y=540 -y -y x=540-y plug it in: 4(540-y)+6y=2340 2160-4y+6y=2340 -2160 -2160 2y=180 divide by 2 on each side y=90

OpenStudy (anonymous):

Any questions?

OpenStudy (livya15):

i worked but you put 540 in the problem instead of 450 :) but i figured it out :)

OpenStudy (anonymous):

oh sorry... :-) good thing you got it!